OpenStudy (anonymous):

how to simplify 8^2/3

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@robtobey can any of you guys help me?

OpenStudy (anonymous):

\[\frac{8^2}{3}=\frac{64}{3} \]

OpenStudy (anonymous):

Sorry for my bad explanation, I mean 8 to the power of 2/3 but thanks anyways.

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

\[8^{2/3}=\sqrt[3]{8^2}=4 \]